# Limits on the Statistical Description of Charged de Sitter Black Holes

### v1

@@ 5 Conclusions
In this work, we investigated the thermodynamics of four-dimensional electrically charged de Sitter black holes, focusing on the ambiguities that arise from the choice of normalization of the timelike Killing vector. We demonstrated that the conventional Gibbons-Hawking normalization leads to an uncommon interpretation of the black hole temperature; it expresses the surface gravity at the horizon as an acceleration measured in the frame of an accelerating observer that is located behind a horizon.
To remedy this, we adopted the Bousso-Hawking normalization of the Killing vector. This choice expresses the different thermodynamic quantities with respect to the frame of the unique freely falling observer at a fixed radial location in between the (outer) black hole horizon and cosmological horizon. This provides a more physically coherent picture, and we derived the corresponding first laws of thermodynamics for both the black hole and cosmological horizon. This led to the introduction of a new mass parameter $\tilde{M}$ whose variation equals the variation of the mass parameter $M$ (defined in the normalization used by Gibbons and Hawking) multiplied by a redshift factor that adapts it to the frame of the freely falling observer.
@@
In contrast to results for the heat capacity using the normalization by Gibbons and Hawking, we find that the heat capacity remains finite and large for Nariai black holes, sufficiently far away from the ultracold point. For cold and ultracold black holes, however, we find the same result for the heat capacity, which vanishes in these extremal limits. Our results therefore suggest that there is no breakdown of the thermodynamic description along the Nariai branch (away from the ultracold point).
For this reason, we also do not expect to see large (quantum) corrections in this limit, and it would be very interesting to confirm this prediction by an explicit one-loop computation that takes into account the physical normalization of the Killing vector. Intriguingly, the results of [38] on Kerr-de Sitter black holes seem to point into this direction. Using the DHS formula [40], the authors of [38] identified whether log-$T$ corrections are present in the different extremal limits from the quasi-normal mode spectrum. Similar to our results, they find log-$T$ corrections for the cold and ultracold branch, but an absence of log-$T$ corrections for Nariai black holes. Although their approach focuses on rotating black holes and does not seem to rely on Bousso-Hawking normalization, we find this connection intriguing and we believe it warrants further investigation. [^10] We believe that their approach effectively implements our choice of normalization through the boundary conditions of the quasinormal modes.
@@
In summary, to fully understand the relation between the different approaches that study corrections to near-extremal black holes in de Sitter space, it is necessary to explicitly compute the one-loop quantum corrections and understand the importance of a physical observer. This perspective seems to be in line with other recent developments that stress the role of the observer in de Sitter space [41, 42, 43, 44, 45, 21].
@@ Acknowledgements
We would like to thank Manus Visser for fruitful discussions at an early stage of this project. LA is supported by the President’s Postdoctoral Fellowship Program at the University of Minnesota and the Heising-Simons Foundation under the “Observational Signatures of Quantum Gravity” collaboration grant 2021-2818. LA thanks the organizers and participants of the workshop “Quantum Gravity in Duluth” where part of this work was presented and acknowledges the University of Wisconsin-Madison for hospitality. PL and GS are supported in part by the DOE grant DE-SC0017647. The work of WS is supported by Starting Grant 2023-03373 from the Swedish Research Council.

### v2

@@ 5 Conclusions
In this work, we investigated the thermodynamics of four-dimensional electrically charged de Sitter black holes, focusing on the ambiguities that arise from the choice of normalization of the timelike Killing vector. We demonstrated explicitly that the conventional Gibbons-Hawking normalization leads to a non-standard interpretation of the black hole temperature; it expresses the surface gravity at the horizon as an acceleration measured in the frame of an accelerating observer that is located behind a horizon.
To remedy this, we adopted the Bousso-Hawking normalization of the Killing vector. This choice expresses the different thermodynamic quantities with respect to the frame of the unique freely falling observer at a fixed radial location in between the (outer) black hole horizon and cosmological horizon. This provides a more physically coherent picture, and we derived the corresponding first laws of thermodynamics for both the black hole and cosmological horizon. This led to the introduction of a new mass parameter $\tilde{M}$ whose variation equals the variation of the mass parameter $M$ (defined in the normalization used by Gibbons and Hawking) multiplied by a redshift factor that adapts it to the frame of the stationary free-falling observer.
@@
In contrast to results for the heat capacity using the normalization by Gibbons and Hawking, we find that the heat capacity remains finite and large for Nariai black holes, sufficiently far away from the ultracold point. For cold and ultracold black holes, however, we find the same result for the heat capacity, which vanishes in these extremal limits. Our results therefore suggest that away from the ultracold point there is no breakdown of the thermodynamic description along the Nariai branch.
For this reason, we also do not expect to see large (quantum) corrections in this limit, and it would be very interesting to confirm this prediction by an explicit one-loop computation that takes into account the physical normalization of the Killing vector. Intriguingly, the results of [42] on Kerr-de Sitter black holes seem to point into this direction. Using the DHS formula [44], the authors of [42] identified whether log-$T$ corrections are present in the different extremal limits from the quasi-normal mode spectrum. Similar to our results, they find log-$T$ corrections for the cold and ultracold branch, but an absence of log-$T$ corrections for Nariai black holes. Although their approach focuses on rotating black holes and does not seem to rely on Bousso-Hawking normalization, we find this connection intriguing and we believe it warrants further investigation [^12] We believe that their approach effectively implements our choice of normalization through the boundary conditions of the quasinormal modes..
@@
In summary, to fully understand the relation between the different approaches that study one-loop corrections to near-extremal black holes in de Sitter space, it is crucial to appreciate and understand their dependence on the selected normalization for the temperature, i.e. the particular physical observer that is being considered. This perspective seems to be in line with other recent developments that stress the role of the observer in de Sitter space [45, 46, 47, 48, 49, 21].
@@ Acknowledgements
We would like to thank Manus Visser for useful discussions. JPvdS would like to thank Zhenbin Yang for an interesting and insightful discussion at the ICTP Workshop on Approaches to Quantum Gravity in de Sitter space. LA is supported by the President’s Postdoctoral Fellowship Program at the University of Minnesota and the Heising-Simons Foundation under the “Observational Signatures of Quantum Gravity” collaboration grant 2021-2818. LA thanks the organizers and participants of the workshop “Quantum Gravity in Duluth” where part of this work was presented and acknowledges the University of Wisconsin-Madison for hospitality. PL and GS are supported in part by the DOE grant DE-SC0017647. The work of WS is supported by Starting Grant 2023-03373 from the Swedish Research Council.
@@ B.2 Charged de Sitter Black Holes
Appendix C Energy in Bousso-Hawking normalization
In the main text, we introduced (the variation of) a new energy parameter $\tilde{M}$. Its definition might seem ad-hoc, so in this Appendix we clarify its interpretation. First, we highlight that from the variational identity
$\delta\tilde{M}=\delta E+P_{\cal O}\delta A_{\cal O}+\tilde{\Phi}_{\cal O}\delta Q~,$ (C.1)
we see that at fixed charge and York boundary, the variation $\delta\tilde{M}$ coincides with the variation of the Brown-York energy.
Second, the variational identity can be integrated to obtain an expression for $\tilde{M}$. To this end, it is useful to write down an ansatz
$\tilde{M}=\frac{M}{\sqrt{f(r_{\cal O})}}+{\cal M}~.$ (C.2)
Varying this ansatz, imposing (2.33), and integrating leads to
${\cal M}=\frac{M}{2f(r_{\cal O})^{3/2}}\left(\int\mathrm{d}M\,\partial_{M}f(r_{\cal O})+\int\mathrm{d}Q\,\partial_{Q}f(r_{\cal O})\right)~.$ (C.3)
Performing these integrals is complicated, but possible in certain specific cases. This gives an explicit expression for $\tilde{M}$.
It should be noted that it is not necessary to introduce $\tilde{M}$ to compute the heat capacity, as we can just use the definition in terms of the heat exchanged ($\tilde{T}\delta S$). We find the energy perspective useful, however, because it highlights that the heat capacity obtained using $\tilde{M}$ coincides with the heat capacity one would have obtained when using the Tolman temperature with a fixed York boundary and fixed charge. This is only true when the York boundary is located at $r=r_{\cal O}$.
This suggests that there are two different perspectives that give the same result for the thermodynamics. One can either absorb all dependence on the York boundary and allow for variations in $r_{\cal O}$ (which is the perspective we took), or impose a York boundary at $r=r_{\cal O}$ and fix its location. We’d like to stress that the equivalence of these two perspectives is only true for our special choice of observer at $r=r_{\cal O}$. Mathematically, the reason is that a variation of $f(r_{\cal O})$ with respect to $r_{\cal O}$ vanishes both when we fix $r_{\cal O}$ (by definition) and also when we do allow for variations because $r_{\cal O}$ is defined to satisfy $f^{\prime}(r_{\cal O})=0$. These observations support the claim that $\tilde{M}$ is the appropriate definition of energy with respect to an observer at $r_{\cal O}$.